Lucy has some money invested at 18% annual interest and three as much invested at 12%. Her total annual income from two investments is 2200pesos. How much was invested at each rate: Lucy has some money invested at 18% annual interest and three as much invested at 12%. Her total annual income from two investments is 2200pesos. How much was invested at each rate
Answer by checkley77(1586)

.18X+.12*3X=2,200
.18X+.36X=2,200
.54X=2,200
X=2,200/.54
X=4,074.07 IS THE AMOUNT INVESTED @ 18%
3*4,074.07= 12,222.21 INVESTED @ 12%.
PROOF:
.18*4,074.07+.12*12,222.21=2,200
733.33+1,466.67=2,200
2,200=2,200